OpenStudy (anonymous):

A2kg iron block is taken from a furnace where its temoerature was 650C and placed on a large block of ice at 0C. Assuming that all the heat given up by the iron is used to melt the ice, how much ice is melted?

OpenStudy (anonymous):

first find the heat that iron is losing to melt ice \[\Large Q=mc \Delta T\] here mass of iron =m=2kg delta t=650-0=650C c=specific heat of iron =442 J/kgC Q= 2*442*650=574600 J divide this 574600 J energy by Latent heat of fusion (because Phase change is taking place) you mill get mass of ICE !! can you find it now ?? let me know if you need further help with this .

OpenStudy (anonymous):

thanks bhai